OpenStudy (anonymous):

Find the equation of the tangent line to the curve y=sin(sin(5x))+e^x^2 at the point where x=0. Your answer must be in the form of an equation for y in terms of x, that is, of the form y=L(x) for some linear function L(x).

OpenStudy (anonymous):

derivative is (cos(cos(5x) 5) +e^x^2 (2x)

hartnn (hartnn):

e^(x^2) or (e^x)^2 ?? these 2 are different

OpenStudy (anonymous):

the first one

hartnn (hartnn):

your derivative of sin sin 5x is incorrect

hartnn (hartnn):

heard about chain rule ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i got the derivative at the top

OpenStudy (anonymous):

i think it's right

hartnn (hartnn):

\((sin (sin 5x))' = cos (sin 5x) \times (sin 5x)' \)

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

my mistake

hartnn (hartnn):

whats your new derivative ?

OpenStudy (anonymous):

cos(sin5x)×(sin5x)′ + e^x^2 (2x)

OpenStudy (anonymous):

cos(sinx5x) * (sin5x)' +e^x^2 2x)

hartnn (hartnn):

(sin 5x)' means derivative of sin 5x which is \( \cos 5x \times 5 \)

hartnn (hartnn):

5cos(sinx5x) * (cos 5x) +2x e^x^2

OpenStudy (anonymous):

okay then now i just have to make it equal to 0?

hartnn (hartnn):

thats y' plug in x=0 to get the slope at that point

OpenStudy (anonymous):

okay?

OpenStudy (anonymous):

then i get the slope?

hartnn (hartnn):

yup

OpenStudy (anonymous):

wouldn't i get 0 for the slope?

hartnn (hartnn):

nope

hartnn (hartnn):

cos 0 =1

OpenStudy (anonymous):

so the slope would be 1?

OpenStudy (anonymous):

no 5

hartnn (hartnn):

bingo! 5 is correct :)

OpenStudy (anonymous):

now what do i do with the 5?

hartnn (hartnn):

you have slope you have x co-ordinate = 0 what else do u need to get the equation of the line ?

hartnn (hartnn):

i mean x co-ordinate of a point

OpenStudy (anonymous):

b?

hartnn (hartnn):

you know you get get the equation of line if you have slope and a point

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

yoriginal - y = m(xoriginal - x) ?

OpenStudy (anonymous):

then solve for y?

hartnn (hartnn):

you need y co-ordinate find y when x=0 plug in x=0 in the original equation

OpenStudy (anonymous):

the answer would be 1 then right?

hartnn (hartnn):

yes so y co-ordinate is 1 point is (0,1) slope is 5 find the equation

OpenStudy (anonymous):

the answer would be y=5x+1

OpenStudy (anonymous):

thanks, again!

hartnn (hartnn):

welcome ^_^